The melting point of bulk gold is 1064 C. In a melting experiment,

it was found that a gold nanocrystal of 5.0 nm in diameter has a

melting point of 817 C, and an unknown gold nanocrystal has a

melting point of 500 C. Estimate the particle size of this unknown

gold sample.
Answer #1

Given, (TMB) melting point of bulk gold = 1064 c = (1064 +273) K = 1337K when Im (817 + 273) K = 1090k, particle diameter, d